OpenStudy (anonymous):

Sergio bought 1 yard of yellow yarn. He used 21 inches of this yarn for a project. What part of the yarn that he bought did he use? Remember, 1 yard equals 36 inches. A.He used 1/36 of the yarn. B.He used 3/10 of the yarn. C.He used 7/12 of the yarn. D.He used 2/3 of the yarn.

OpenStudy (hoblos):

21/36 = 7/12

OpenStudy (anonymous):

\[21\div36 = 7/12\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can somebody actually show me how they got it